Harvey Industries Pays a current dividend of $6.10 and shareholders require a 12% return. The dividend will grow at a high rate of 20% and then gradually decline to 5% over a six-year period. The value of Harvey Industries shares using the H Model is closest to:

a. $121.39.

b. $127.74.

c. $130.71.

d. $137.93.
Answer #1

Formula to compute stock value using H-model is:

Stock value = [D0(1+g2) + D0 x H x (g1 – g2)]/ (r – g2)

D0 = The most recent dividend payment = $ 6.10

g1 = The initial high growth rate = 20 % or 0.2

g2 = The terminal growth rate = 5 % or 0.05

r = The discount rate = 12 % or 0.12

H = The half-life of high growth period = 6 /2 = 3 years

Substituting all the values on above formula, we get stock value as:

Stock value = [$ 6.10(1+0.05) + $ 6.10 x 3 x (0.2 – 0.05)]/ (0.12 – 0.05)

                    = [($ 6.10 x 1.05) + ($ 6.10 x 3 x 0.15)]/ 0.07

                    = ($ 6.405 + $ 2.745) / 0.07

                     = $ 9.15/ 0.07 = $ 130.714285714286 or $ 130.71

The value of Harvey Industries shares using H-model is closest to $ 130.71

Hence option “c. $ 130.71” is correct answer.